OpenStudy (anonymous):

Write the inverse variation equation of the following. Y varies inversely with x and y = -12 when x = 6. A. xy = -2 B. xy = -72 C. xy = -0.5

OpenStudy (lgbasallote):

is Y different from y?

OpenStudy (anonymous):

umm...it's possible

OpenStudy (anonymous):

y=k/x yx=k

OpenStudy (lgbasallote):

ill just assume they're equal so the first equation is y = k/x -12 = k/6 k = -72 so y = k/x y = -72/x xy = -72

OpenStudy (anonymous):

oh okay i think i get how to do it now...thnk u igba and shadowone.